Directions: Solve the equation by completing the square. Express any complex numbers using i notation.

2x^2+4x+1=0

Answer by stanbon(75887)

2x^2+4x+1=0
Divide thru by 2 to get:
x^2+2x = -1/2
Complete the square on the left side, as follows:
x^2+2x+1 = -1/2 + 1
(x+1)^2 = +1/2
x+1=(sqrt2)/2 or x+1 = -(sqrt2)/2
x=-1+(sqrt2)/2 or x=-1-(sqrt2)/2
Cheers,
Stan H.